OpenStudy (anonymous):

Find the zeros of f. f(x)= x^7 (4e^(4x)) + 7x^6e^(4x)

myininaya (myininaya):

\[x^7*4e^{4x}+7x^6e^{4x}=x^6e^{4x}(4x+7)=0\] so we have x=0 and x=-7/4

OpenStudy (anonymous):

Ohhhhhhh I see, you factor it out!!! Thanks!!!!